OpenStudy (anonymous):

The hiking club wants to sell 50 pounds of an energy snack food that will cost them only $1.80 per pound. The mixture will be made up of nuts that cost $2.00 per pound and raisins that cost $1.50 per pound. How much of each should they buy to obtain 50 pounds of the desired mixture?

OpenStudy (slaaibak):

say n = pounds of nuts r = pounds or raisins so, the first constraint is r + n = 50 If it must cost 1.8 per pound, 50 pounds will cost 50 * 1.8 = 90 so n * 2 + 1.5 * r = 90 Now the two equations are: n + r = 50 2n + 1.5r = 90 Solve that to get the answer